# Competitive Equilibrium for Chores: from Dual Eisenberg-Gale to a Fast, Greedy, LP-based Algorithm

### v1

@@ 2. Preliminaries
In a CE in the chores market, we determine price $p_{j}$ for each chore $j$, and an allocation $x\in\mathbb{R}^{n\times m}_{\geq 0}$ of chores to the agents, where each $x_{ij}$ is the amount of chore $j$ allocated to agent $i$, such that (i) each agent $i$ gets a disutility minimizing bundle $x_{i}=(x_{i1},x_{i2},\dots,x_{im})$, subject to her earning requirement of $B_{i}$ units of money, and (ii) all chores are allocated. Formally,
Definition 1 (Competitive Equilibrium for Chores).
@@
Condition (E1) and (E2) ensure that each agent gets a disutility-minimizing bundle subject to their earning constraint, and (E3) captures that all chores are allocated in this process. CEEI is one of the most well-studied special cases of CE, where $B_{i}=1$, for all $i\in[n]$. It is a well known [CITE] that if $(p,x)$ satisfy CEEI, then (i) $x$ is envy-free, i.e., $d_{i}(x_{i})\leq d_{i}(x_{j})$ for all $i,j\in[n]$, and (ii) $x$ is Pareto-optimal, i.e., there is no $y$ such that $d_{i}(y_{i})\leq d_{i}(x_{i})$ with at least one strict inequality. CEEI satisfies many other desirable properties like core-stability, and we refer the reader to (Bogomolnaia et al., 2017) for a more detailed explanation of the remarkable fairness and efficiency properties satisfied by CEEI. The existence of CE in the chores market can be shown by adapting the fixed point arguments in (Arrow and Debreu, 1954)(Bogomolnaia et al., 2017) show an interesting characterization of all CE in the chores market. They show that there is a one-to-one correspondence between the set of CE and all KKT points of Program 1.
$\inf_{x\geq 0}\;$ $\prod_{i\in[n]}\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}^{B_{i}}$
(1) s.t. $\sum_{i\in[n]}x_{ij}=1\quad\forall j\in[m]$
$\sum_{j\in[m]}d_{ij}x_{ij}>0\quad\forall i\in[n].$
Observe crucially that the open constraint $\sum_{j\in[m]}d_{ij}x_{ij}>0\quad\forall i\in[n]$ rules out allocations where the one agent has zero disutility, and the global infimum is achieved at the closure of these constraints. In this paper, we work with an equivalent formulation (Program 2), where we change our objective from the product of disutilties to the sum of logarithms of disutilties.
(2) $\inf_{x\geq 0}\;$ $\sum_{i\in[n]}B_{i}\cdot\log\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}$
s.t. $\sum_{i\in[n]}x_{ij}=1\quad\forall j\in[m].$
Open constraint to Unbounded Objective. While the domain of the log function dictates that the total disutility of every agent is strictly positive [^9] Therefore, we do not need to explicitly impose the open constraints., the program admits unbounded-objective directions – the objective tends to negative infinity as the disutility of an agent tends to zero. Furthermore, first and second order methods are attracted towards this direction (Boodaghians et al., 2022).
Computing CE in Chores Market.
The exact complexity of finding CEEI in the chores market is open. However, there are algorithms (Boodaghians et al., 2022; Chaudhury et al., 2022) that compute an approximate CEEI in the chores market in polynomial time. Intuitively, in an approximate CE, conditions (E1), (E2), and (E3) in Definition 1 are satisfied approximately. Formally,
Definition 2 ($\epsilon$-approximate CE (Boodaghians et al., 2022)).
@@
Bregman divergence and generalized strong convexity.
@@
Let $h:X\rightarrow\mathbb{R}$ be a differentiable, strictly convex function defined on a convex set $X$. The Bregman divergence associated with $h$ between $y,x\in X$ is defined as $D_{h}(y\,\|\,x)=h(y)-h(x)-\langle\nabla h(x),y-x\rangle$. $h$ is called the reference function of $D_{h}(\cdot\,\|\,\cdot)$. For example, if $h(x)=\frac{1}{2}{\left\|x\right\|}^{2}$ then $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$. Since $h$ is convex, $D_{h}(y\,\|\,x)\geq 0$ for all $y,x\in X$.
@@
We say a differentiable function $f$ is $\alpha$-strongly convex with respect to $h$ if
(3) $\text{dom}\,f\subseteq X\quad\text{and}\quad f(y)\geq f(x)+\langle\nabla f,y-x\rangle+\alpha D_{h}(y\,\|\,x)\text{ for any }y,x\in\text{dom}\,f.$
@@
We refer Eq. 3 with an arbitray $D_{h}(\cdot\,\|\,\cdot)$ as generalized strong convexity or relatively strong convexity, and it reduces to (standard) strong convexity if $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$.
3. The Convex Maximization Program
In this Section, we introduce a novel program with no unbounded-objective directions, such that there is a one-to-one correspondence between CE in the chores market and the KKT points of our program. Our inspiration for this program comes from the dual program for CE in the goods setting. In the context of goods, the Eisenberg-Gale convex program has a dual program as follows (Cole et al., 2017)
$\min_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$
s.t. $p_{j}\geq\beta_{i}v_{ij}\quad\forall i,j.$
@@ 3. The Convex Maximization Program
This program has been used algorithmically in several ways. The proportional response dynamics can be derived by a change of variable and dualization step on the “goods” EG dual (Zhang, 2011; Birnbaum et al., 2011), and the PACE algorithm for online Fisher market equilibrium is also derived from this program (Gao et al., 2021; Liao et al., 2022). Given the algorithmic usefulness of the EG dual in the goods case, it is natural to ask whether there is an analogue for the chores case. The EG program for chores, however, is not a convex program, and thus convex duality does not let us directly derive a dual. Nonetheless, we may guess that the chores EG “dual” should be the maximization version of the goods EG dual, analogously to what occurs in the primal space. This guess yields the following convex maximization program
(Chores Dual) $\sup_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$
s.t. $p_{j}\leq\beta_{i}d_{ij}\quad\forall i,j.$
@@
Henceforth we shall refer to this as the chores dual, even though it is not a dual program in the formal sense of convex duality. It is worth noting that, for example, we do not have global weak duality: some feasible $(\beta,p)$ pairs to Chores Dual yield a dual objective that does not lower bound the primal objective (this is easy to see, since the dual objective tends to infinity along some rays).
@@
Theorem 1 ().
There is a one-to-one correspondence between competitive equilibria of a chores Fisher market and the KKT points of Chores Dual.
@@
Let $x$ be the Lagrange multiplier variables for the constraint in Chores Dual. By the KKT conditions, $(p,\beta,x)$ is a KKT point of Chores Dual if and only if it satisfies
@@
Note the $\beta>0$ (instead of $\beta\geq 0$) in $(i)$ above. This is implied by the domain constraint of the log function on $\beta_{i}$s.
First, we show that any KKT point of Chores Dual is a CE. By (ii), we have $\frac{p_{j}}{d_{ij}}=\beta_{i}\geq\frac{p_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$, if $x_{ij}>0$ for all $i\in[n],j\in[m]$, which is equivalent to $(E2)$ in Definition 1. Combining (ii) and (iv), we have $\sum_{j}p_{j}x_{ij}=\sum_{j}d_{ij}\beta_{i}x_{ij}=B_{i}$ for all $i\in[n]$. Also, (iii) leads to $\sum_{i}x_{ij}=1$ if $p_{j}>0$ for all $j\in[m]$.
Next, we prove that any CE is a KKT point of Chores Dual. Given any CE $(p,x)$, we denote $\beta_{i}:=\frac{B_{i}}{\sum_{j}d_{ij}x_{ij}}$ for all $i\in[n]$. Then, $(p,\beta,x)$ trivially satisfies (i), (iii) and (iv). Since, for all $i$, $\sum_{j}d_{ij}x_{ij}=\max_{x^{\prime}_{i}:\langle p,x^{\prime}_{i}\rangle\geq B_{i}}\sum_{j}d_{ij}x^{\prime}_{ij}$ and $\sum_{j}d_{i}jx_{i}j\leq\frac{d_{ij^{\prime}}}{p_{j}}\cdot B_{i}$ for all $j^{\prime}$, we have $d_{ij}\beta_{i}=\frac{d_{ij}B_{i}}{\sum_{j}d_{ij}x_{ij}}\geq\frac{d_{ij}B_{i}}{d_{ij}({B_{i}}/{p_{j}})}=p_{j}$ for all $i\in[n],j\in[m]$. For all $i\in[n]$, $\sum_{j\in[m]}x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)=B_{i}-B_{i}=0$ since $\langle p,x_{i}\rangle=B_{i}$. Then, (ii) follows since $x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)\leq 0$ for all $i\in[n],j\in[m]$. ∎
Next we look at another property typically satisfied by primal and dual convex programs: strong duality implies that optimal primal and dual solutions have the same objective. As we already noted, our Chores Dual does not even satisfy weak duality. Yet, it turns out that we do have a local form of strong duality, in the sense that the objectives are equal at KKT points. To see this, we next write more explicit versions of the primal and dual, in a way that keeps various important constants around for showing this equality (these same constants are needed in the goods setting).
@@
(Chores Primal) $\inf_{x\geq 0}\;$ $\sum_{i}B_{i}\log(u_{i})$
s.t. $\langle d_{i},x_{i}\rangle\leq u_{i}\quad\forall\;i\in[n]$ $\quad[\beta_{i}]$
$\sum_{i}x_{ij}\geq 1\quad\forall\;j\in[m]$ $\hskip 20.0pt[p_{j}]$
@@
(Chores Dual’) $\sup_{p\geq 0}\;$ $\sum_{j}p_{j}-\sum_{i}\Big{(}B_{i}\log(\beta_{i})+B_{i}\log{B_{i}}-B_{i}\Big{)}$
s.t. $p_{j}\leq d_{ij}\beta_{i}\quad\forall\;i\in[n],j\in[m]$ $[x_{ij}].$
Theorem 2 ().
If $(x,u,\beta,p)$ is a KKT point of Chores Primal, then $(p,\beta,x)$ is a KKT point of Chores Dual’, and vice versa. Furthermore, the primal and dual program objectives are equal at every KKT point.
@@
The first part is obvious due to the one-to-one correspondence between KKT points and CE for both programs. Since $(x,u,\beta,p)$ is a KKT point of Chores Primal, we have $\beta_{i}=B_{i}/u_{i}$ for all $i\in[n]$. Since $(p,\beta,x)$ is a KKT point of Chores Dual’, and thus a CE of the corresponding chores Fisher market, we have $\sum_{j}p_{j}=\sum_{j}\sum_{i}p_{j}x_{ij}=\sum_{i}\sum_{j}p_{j}x_{ij}=\sum_{i}B_{i}$. By these, we can obtain that
$\sum_{j}p_{j}-\sum_{i}\Big{(}B_{i}\log(\beta_{i})+B_{i}\log{B_{i}}-B_{i}\Big{)}=\sum_{i}B_{i}\log{\frac{B_{i}}{\beta_{i}}}+\sum_{j}p_{j}-\sum_{i}B_{i}=\sum_{i}B_{i}\log{u_{i}}.$
Similarly to the primal EG for chores, the dual EG for chores has problems from an optimization perspective. One might try to find a KKT point by finding a local maximum of Chores Dual using some sort of iterative optimization method on the interior of the convex feasible region. However, this runs into trouble because Chores Dual has many directions where the objective tends to infinity, generally because we can either let $\beta_{i}$ tend to zero in order for $-\log\beta_{i}$ to tend to infinity, or by letting the prices tend to infinity, which grows faster than $-\log\beta_{i}$.
Next, we show what we found to be a highly surprising fact: the Chores Dual can be “fixed” such that all these directions that lead to an unbounded objective are avoided, without removing any KKT points. This is achieved by adding a new constraint, which is redundant from a CE perspective. It is easy to see that in any CE, we must have that every buyer earns their budget exactly (otherwise they could do fewer chores and still satisfy their earning constraint), and since every item is allocated exactly, we must have that $\sum_{j}p_{j}=\sum_{i}B_{i}$. Adding this constraint to Chores Dual yields the following convex maximization program:
(Chores Dual Redundant) $\max_{\beta\geq 0,p\geq 0}\quad$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$
@@
The new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ in Chores Dual Redundant cuts off directions with infinite objective: 1) prices are upper bounded by $\sum_{i}B_{i}$, so the rays where prices go to infinity are avoided, 2) each $\beta_{i}$ is lower bounded by $\max_{j}p_{j}/d_{ij}\geq\sum_{i}B_{i}/(m\cdot\max_{j}d_{ij})$ since $p_{j}\geq\sum_{i}B_{i}/m$ for some chore $j$. The following lemma formalizes this fact; the proof follows from the above discussion.
Lemma 3 ().
The optimal value of Chores Dual Redundant is at most
$\left(\sum_{i}B_{i}\right)\left(1-\log(\sum_{i}B_{i})+\log(m)+\log(\max_{i,j}d_{ij})\right).$
Crucially, it turns out that the new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ is redundant, in the sense that its Lagrange multiplier is zero at every KKT point. This implies that the one-to-one correspondence between CE and KKT points still holds for Chores Dual Redundant.
Theorem 4 ().
There is a one-to-one correspondence between CE of the chores Fisher market and the KKT points of Chores Dual Redundant.
@@
Here, we refer to the two constraints in Chores Dual Redundant as the optimal disutility constraints (the first constraints) and the redundant constraint (the second constraint). Let $x$ and $\mu$ be the dual variables corresponding to the optimal disutility constraint and redundant constraint, respectively. By the KKT conditions, $(p,\beta,x,\mu)$ is a KKT point of Chores Dual Redundant if and only if it satisfies the conditions (i), (ii) and (iv) in Theorem 1as well as
@@
    $\sum_{i}x_{ij}+\mu\geq 1$ and $p_{j}\left(\sum_{i}x_{ij}+\mu-1\right)=0$ for all $j\in[m]$
First, we show that any KKT point of Chores Dual Redundant is a CE. To do this, we only need to show $\mu=0$ if $(p,\beta,x,\mu)$ is a KKT point of Chores Dual Redundant, since the conditions (ii) and (iv) guarantee that every agent chooses her optimal bundle.
@@
Next, we prove that any CE is a KKT point of Chores Dual Redundant. Set $\mu=0$ and the rest follows similarly to the second part of the proof of Theorem 1. ∎
@@
Because Chores Dual Redundant does not have unbounded-objective directions (Lemma 3), it allows the use of first-order methods and interior-point methods on the convex polytope.
$Y=\big{\{}(\beta,p):p_{j}\leq\beta_{i}d_{ij},\sum_{j}p_{j}=\sum_{i}B_{i}\big{\}}.$

### v2

@@ 2.1 Fisher Market with Chores
In a CE for a chores market, we wish to find a price $p_{j}$ for each chore $j$, and an allocation $x\in\mathbb{R}^{n\times m}_{\geq 0}$ of chores to the agents, where each $x_{ij}$ is the amount of chore $j$ allocated to agent $i$, such that (i) each agent $i$ gets a disutility minimizing bundle $x_{i}=(x_{i1},x_{i2},\dots,x_{im})$, subject to her earning requirement of $B_{i}$ units of money, and (ii) all chores are allocated. Formally,
Definition 1.
@@
Condition (E1) and (E2) ensure that each agent gets a disutility-minimizing bundle subject to her earning constraint, and (E3) captures that all chores are allocated in this process. CEEI is one of the most well-studied special cases of CE, where $B_{i}=1$, for all $i\in[n]$. It is a well known (see Bogomolnaia et al. (2017, Lemma 1)) that if $(p,x)$ satisfy CEEI, then (i) $x$ is envy-free, i.e., $d_{i}(x_{i})\leq d_{i}(x_{j})$ for all $i,j\in[n]$, and (ii) $x$ is Pareto-optimal, i.e., there is no $y$ such that $d_{i}(y_{i})\leq d_{i}(x_{i})$ with at least one strict inequality. CEEI satisfies many other desirable properties like core-stability, and we refer the reader to Bogomolnaia et al. (2017) for a more detailed explanation of the remarkable fairness and efficiency properties satisfied by CEEI. The existence of CE in the chores market can be shown by adapting the fixed point arguments in Arrow and Debreu (1954).
Our algorithms will be concerned with computing approximate CE solutions. To that end, we relax conditions (E1), (E2), and (E3) in Definition 1 as follows,
Definition 2 ($\epsilon$-approximate CE (Boodaghians et al. 2022)).
@@
2.2 Mathematical Programming Approaches to Competitive Equilibria
In the case of CE for goods, it is known that a CE allocation can be obtained from the Eisenberg-Gale convex program. In the goods case with linear utilities, each buyer has a value vector $v_{i}\in\mathbb{R}_{+}^{m}$ describing their value per good, and they compute demands by maximizing their obtained value subject to budget constraints. The EG program for this case is stated on the left below.
[…]
On the right in (1) is the dual of the EG program, which minimizes a function of $p$ and $\beta$, where $p$ and $\beta$ can be interpreted as prices for the goods, and inverse bang-per-buck variables for the buyers, respectively. We will use the dual of the goods EG program as inspiration for our chores approach.
Bogomolnaia et al. (2017) show an interesting characterization of all CE in the chores market. Similar to the case of goods, they show that for chores there is also an EG-style program for computing a CE. More specifically, they show that there is a one-to-one correspondence between the set of CE and all KKT points of the following program,
$\inf_{x\geq 0}$ $\prod_{i\in[n]}\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}^{B_{i}}$ (2)
s.t. $\sum_{i\in[n]}x_{ij}=1\quad\;\;\;\;\;\forall\,j\in[m]$
$\sum_{j\in[m]}d_{ij}x_{ij}>0\quad\forall\,i\in[n].$
Applying the logarithm, which is a monotone transformation, to the objective yields a mathematical program analogous to the EG program for goods,
$\inf_{x\geq 0}$ $\sum_{i\in[n]}B_{i}\cdot\log\bigg{(}\sum_{j\in[m]}d_{ij}x_{ij}\bigg{)}$ (3)
s.t. $\sum_{i\in[n]}x_{ij}=1\quad\;\;\;\;\;\forall\,j\in[m].$
After applying the logarithm, we no longer need an explicit open constraint, since the domain constraint on the logarithm already enforces strictly positive disutilities.
However, this program has several disadvantages compared to the goods case. Most notably, it is a non-convex program, and the objective function can tend to negative infinity within the feasible region (as the disutility of an agent tends to zero). This makes it challenging algorithmically, because iterative methods that operate inside the feasible set are attracted to these “negatively-infinite-objective” points (Boodaghians et al. 2022).
2.3 Bregman Divergence and Generalized Strong Convexity
@@
Let $h:\Omega\rightarrow\mathbb{R}$ be a differentiable, strictly convex function defined on a convex set $\Omega$. The Bregman divergence associated with $h$ between $y,x\in\Omega$ is defined as $D_{h}(y\,\|\,x)=h(y)-h(x)-\langle\nabla h(x),y-x\rangle$. The function $h$ is called the reference function of $D_{h}(\cdot\,\|\,\cdot)$. For example, if $h(x)=\frac{1}{2}{\left\|x\right\|}^{2}$ then $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$. Since $h$ is strictly convex, $D_{h}(y\,\|\,x)\geq 0$ for all $y,x\in\Omega$, and $D_{h}(y\,\|\,x)=0$ iff $y=x$.
@@
We say a differentiable function $f$ is $\alpha$-strongly convex with respect to $h$ if $\text{dom}\,f\subseteq\Omega$ and
$f(y)\geq f(x)+\langle\nabla f(x),y-x\rangle+\alpha D_{h}(y\,\|\,x)\quad\forall\,y,x\in\text{dom}\,f.$ (4)
@@
We refer Eq. 4 with an arbitrary $D_{h}(\cdot\,\|\,\cdot)$ as generalized strong convexity or relatively strong convexity, and it reduces to (standard) strong convexity if $D_{h}(y\,\|\,x)=\frac{1}{2}{\left\lVert y-x\right\|}^{2}$.
3 The Chores Dual Program
@@
In this Section, we introduce a novel program with no poles, such that there is a one-to-one correspondence between CE in the chores market and the KKT points of our program. Our inspiration for this program comes from the dual program for CE in the goods setting (Eq. 1). In the goods case, the dual program has been used extensively in order derive various algorithmic approaches for computing CE. The proportional response dynamics can be derived by a change of variable and dualization step on the goods EG dual (Zhang 2011, Birnbaum et al. 2011), and the PACE algorithm for online Fisher market equilibrium is also derived from this program (Gao et al. 2021, Liao et al. 2022). Given the algorithmic usefulness of the EG dual in the goods case, it is natural to ask whether there is an analogue for the chores case. The EG program for chores, however, is not a convex program, and thus convex duality does not let us directly derive a dual. Nonetheless, we may guess that the chores EG “dual” should be the maximization version of the goods EG dual, analogously to how the chores “primal” is the minimization version of the goods EG primal. This guess yields the following program which maximizes a convex function:
$\sup_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$ (Chores Dual)
s.t. $\;p_{j}\leq\beta_{i}d_{ij}\quad\quad\forall\;i,j.$
@@
Henceforth we shall refer to this as the chores dual, even though it is not a dual program in the formal sense of convex duality. It is worth noting that, for example, we do not have global weak duality: some feasible pairs $(\beta,p)$ of (Chores Dual) yield a dual objective that does not lower bound the primal objective (this is easy to see, since the dual objective tends to infinity near poles).
@@
Theorem 1.
There is a one-to-one correspondence between competitive equilibria of a chores Fisher market and the KKT points of (Chores Dual).
@@
First, we show that any KKT point of (Chores Dual) is a CE. Let $(\tilde{p},\tilde{\beta})$ denote a feasible point in the primal space of (Chores Dual), and $\tilde{x}$ be the dual variable associated with the constraint in (Chores Dual). Note that we overload notation slightly here: $(p,\beta,x)$ can, at this stage, only be understood to be primal and dual variables that arise from the programs, whereas we previously used these notations to refer to market quantities like prices and allocations. We will show that the primal and dual variables do in fact have these interpretations, thus justifying the notation. By the KKT conditions, any KKT point $(\tilde{p},\tilde{\beta},\tilde{x})$ of (Chores Dual) satisfies
@@
Note the $\beta>0$ (instead of $\beta\geq 0$) in $(i)$ above. This is implied by the domain constraint of the log function on $\beta_{i}$s. By (ii), we have $\frac{\tilde{p}_{j}}{d_{ij}}=\tilde{\beta}_{i}\geq\frac{\tilde{p}_{j^{\prime}}}{d_{ij^{\prime}}}$ for all $j^{\prime}\in[m]$, if $\tilde{x}_{ij}>0$ for all $i\in[n],j\in[m]$, which is equivalent to $(E2)$ in Definition 1. Combining (ii) and (iv), we have $\sum_{j}\tilde{p}_{j}\tilde{x}_{ij}=\sum_{j}d_{ij}\tilde{\beta}_{i}\tilde{x}_{ij}=B_{i}$ for all $i\in[n]$. Also, (iii) leads to $\sum_{i}\tilde{x}_{ij}=1$ if $\tilde{p}_{j}>0$ for all $j\in[m]$. Thus, by definition, $(\tilde{p},\tilde{x})$ constitutes a CE if we regard $p$ as price and $x$ as allocation.
Next, we prove that any CE is a KKT point of (Chores Dual). Given any CE $(p,x)$, we denote $\beta_{i}:=\frac{B_{i}}{\sum_{j}d_{ij}x_{ij}}$ for all $i\in[n]$. Then, $(p,\beta,x)$ trivially satisfies (i), (iii) and (iv). Since, for all $i$, $\sum_{j}d_{ij}x_{ij}=\max_{x^{\prime}_{i}:\langle p,x^{\prime}_{i}\rangle\geq B_{i}}\sum_{j}d_{ij}x^{\prime}_{ij}$ and $\sum_{j}d_{i}jx_{i}j\leq\frac{d_{ij^{\prime}}}{p_{j}}\cdot B_{i}$ for all $j^{\prime}$, we have $d_{ij}\beta_{i}=\frac{d_{ij}B_{i}}{\sum_{j}d_{ij}x_{ij}}\geq\frac{d_{ij}B_{i}}{d_{ij}({B_{i}}/{p_{j}})}=p_{j}$ for all $i\in[n],j\in[m]$. For all $i\in[n]$, $\sum_{j\in[m]}x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)=B_{i}-B_{i}=0$ since $\langle p,x_{i}\rangle=B_{i}$. Then, (ii) follows since $x_{ij}\left(p_{j}-d_{ij}\beta_{i}\right)\leq 0$ for all $i\in[n],j\in[m]$. Hence, $(p,\beta,x)$ satisfies KKT conditions. ∎
Next we look at another property typically satisfied by primal and dual convex programs: strong duality implies that optimal primal and dual solutions have the same objective. As we already noted, our (Chores Dual) does not even satisfy weak duality. Yet, it turns out that we do have a local form of strong duality, in the sense that the objectives are equal at KKT points. To see this, we next write more explicit versions of the primal and dual, in a way that keeps various important constants around for showing this equality (these same constants are needed in the goods setting).
@@
$\inf_{x\geq 0,u\geq 0}$ $f^{p}(x,u):=\sum_{i}B_{i}\log(u_{i})$ (5)
s.t. $\langle d_{i},x_{i}\rangle\leq u_{i}\quad\hskip 20.0pt\forall\;i\in[n]$ $\quad(\beta_{i})$
$-\sum_{i}x_{ij}\leq-1\quad\;\forall\;j\in[m]$ $\hskip 20.0pt(p_{j})$
@@
$\sup_{\beta\geq 0,p\geq 0}$ $f^{d}(\beta,p):=\sum_{j}p_{j}-\sum_{i}B_{i}\log{\beta_{i}}+\sum_{i}\big{(}B_{i}\log{B_{i}}-B_{i}\big{)}$ (6)
s.t. $p_{j}\leq d_{ij}\beta_{i}\quad\quad\quad\forall\;i\in[n],\,j\in[m]\hskip 30.0pt(x_{ij}).$
Theorem 2.
If $(x,u,\beta,p)$ is a KKT point [^4] Any pair of primal and dual variables which satisfies the KKT conditions is called a KKT point. of (5), then $(\beta,p,x)$ is a KKT point of (6). Reversely, if $(\beta,p,x)$ is a KKT point of (6) and $u_{i}=\langle d_{i},x_{i}\rangle$ for all $i\in[n]$, then $(x,u,\beta,p)$ is a KKT point of (5). Furthermore, if $(x,u,\beta,p)$ is a KKT point of (5), then
$f^{p}(x,u)=f^{d}(\beta,p).$
@@
The first part is obvious due to the one-to-one correspondence between KKT points and CE for both programs. Since $(x,u,\beta,p)$ is a KKT point of (5), we have $\beta_{i}=\frac{B_{i}}{u_{i}}$ for all $i\in[n]$. Since $(p,\beta,x)$ is a KKT point of (6), and thus a CE of the corresponding chores Fisher market, we have $\sum_{j}p_{j}=\sum_{j}\sum_{i}p_{j}x_{ij}=\sum_{i}\sum_{j}p_{j}x_{ij}=\sum_{i}B_{i}$. By these, we can obtain that
$\sum_{j}p_{j}-\sum_{i}B_{i}\log{\beta_{i}}+\sum_{i}\big{(}B_{i}\log{B_{i}}-B_{i}\big{)}=\sum_{i}B_{i}\log{\frac{B_{i}}{\beta_{i}}}+\sum_{j}p_{j}-\sum_{i}B_{i}=\sum_{i}B_{i}\log{u_{i}}.$
Similarly to the primal EG for chores, the dual EG for chores has problems from an optimization perspective. One might try to find a KKT point by finding a local maximum of (Chores Dual) using some sort of iterative optimization method on the interior of the convex feasible region. However, this runs into trouble because (Chores Dual) has two ways that the objective tends to infinity, generally because we can either let $\beta_{i}$ tend to zero in order for $-\log\beta_{i}$ to tend to infinity (we refer to these as poles, as in the primal case), or by letting the prices tend to infinity, since their contribution to the objective grows faster than $-\log\beta_{i}$.
In Fig. 1(a), we show a diagram which illustrates how naive gradient descent methods fail because they are either attracted to poles, or move along an unbounded direction. This instance is a Fisher market with $2$ agents and $1$ chore. In this instance, $B_{1}=B_{2}=1$, $d_{11}=2$ and $d_{21}=1$. Obviously, there is only one KKT point / CE in this instance: $p_{1}=2$; $\beta_{1}=1,\beta_{2}=2$. As we can see, even starting from some feasible points that are very close to the KKT point, trajectories generated by the gradient ascending algorithm are attracted by either of those poles.
Figure 1: An instance of a Fisher market with $2$ agents and $1$ chore. The gray region (polyhedral) is the feasible region of (Chores Dual). The (only) KKT point / CE ($(\beta_{1},\beta_{2},p_{1})=(1,2,2)$) is marked by the yellow point. The blue arrow denotes the direction of gradient at the KKT point. In the left, two green trajectories are iterates generated by the gradient ascending algorithm, starting from two feasible points. In the right, the blue region (2-dimensional polyhedral) is the feasible region of (Chores Dual Redundant).
Next, we show what we found to be a highly surprising fact: the (Chores Dual) can be “fixed” such that both poles and unbounded directions are avoided, without removing any KKT points. This is achieved by adding a new constraint, which is redundant from a CE perspective. It is easy to see that in any CE, we must have that every buyer earns their budget exactly (otherwise they could do fewer chores and still satisfy their earning constraint), and since every item is allocated exactly, we must have that $\sum_{j}p_{j}=\sum_{i}B_{i}$. Adding this constraint to (Chores Dual) yields the following program:
$\max_{\beta\geq 0,p\geq 0}$ $\sum_{j}p_{j}-\sum_{i}B_{i}\log(\beta_{i})$ (Chores Dual Redundant)
@@
The new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ in (Chores Dual Redundant) cuts off directions with infinite objective: 1) prices are upper bounded by $\sum_{i}B_{i}$, so the rays where prices go to infinity are avoided, 2) each $\beta_{i}$ is lower bounded by $\max_{j}p_{j}/d_{ij}\geq\sum_{i}B_{i}/(m\cdot\max_{j}d_{ij})$ since $p_{j}\geq\sum_{i}B_{i}/m$ for some chore $j$. The following lemma formalizes this fact; the proof follows from the above discussion.
Lemma 1.
The optimal value of (Chores Dual Redundant) is at most
$\Big{(}\sum_{i}B_{i}\Big{)}\Big{(}1-\log(\sum_{i}B_{i})+\log(m\max_{i,j}d_{ij})\Big{)}.$
Crucially, it turns out that the new constraint $\sum_{j}p_{j}=\sum_{i}B_{i}$ is redundant, in the sense that its Lagrange multiplier is zero at every KKT point. The redundancy also means that this constraint is implied by the KKT conditions of (Chores Dual), which implies that the one-to-one correspondence between CE and KKT points still holds for (Chores Dual Redundant).
Theorem 3.
There is a one-to-one correspondence between CE of the chores Fisher market and the KKT points of (Chores Dual Redundant).
@@
Here, we refer to the two constraints in (Chores Dual Redundant) as the optimal disutility constraints (the first constraints) and the redundant constraint (the second constraint). Let $x$ and $\mu$ be the dual variables corresponding to the optimal disutility constraint and redundant constraint, respectively. By the KKT conditions, $(p,\beta,x,\mu)$ is a KKT point of (Chores Dual Redundant) if and only if it satisfies the conditions (i), (ii) and (iv) in Theorem 1 as well as
@@
    $\sum_{i}x_{ij}+\mu\geq 1$ and $p_{j}\left(\sum_{i}x_{ij}+\mu-1\right)=0$ for all $j\in[m]$.
First, we show that any KKT point of (Chores Dual Redundant) is a CE. To do this, we only need to show $\mu=0$ if $(p,\beta,x,\mu)$ is a KKT point of (Chores Dual Redundant), since the conditions (ii) and (iv) guarantee that every agent chooses her optimal bundle.
@@
Next, we prove that any CE is a KKT point of (Chores Dual Redundant). Set $\mu=0$ and the rest follows similarly to the second part of the proof of Theorem 1. ∎
In Fig. 1(b) we show how the redundant constraint restricts the feasible set in the instance in Fig. 1(a) in a way that avoids poles.
@@
Because (Chores Dual Redundant) does not have poles (Lemma 1), it allows the use of first-order methods and interior-point methods on the convex polyhedron
$Y=\Big{\{}(\beta,p)\in\mathbb{R}^{n+m}_{\geq 0}:\,p_{j}\leq\beta_{i}d_{ij},\;\forall\,i\in[n],j\in[m],\;\sum_{j\in[m]}p_{j}=\sum_{i\in[n]}B_{i}\Big{\}}.$